A student followed the procedure in this experimen
Answer #1

Given,

The volume of undiluted juice = 20.0 mL x ( 1L /1000 mL) = 0.02 L

The concentration of NaOH = 9.580 x 10-2 M

The volume of NaOH solution required for the titration = 12.84 mL x ( 1L /1000 mL) = 0.01284 L

1) Calculating the number of moles of NaOH required for the titration,

We know, the formula,

Molarity = Number of moles / L of solution

9.580 x 10-2 M = Number of moles / 0.01284 L

Number of moles = 1.23 x 10-3 mol of NaOH

2) Calculating the number of moles of C3H5O(COOH)3 titrated,

The reaction between acid and NaOH is,

C3H5O(COOH)3(aq) + 3NaOH(aq) \rightarrow C3H5O(COONa)3(aq) + 3H2O(l)

Using the calculated number of moles of NaOH and the mole ratio from the balanced chemical reaction, calculating the number of moles of C3H5O(COOH)3,

= 0.00123 mol of NaOH x ( 1 mol of C3H5O(COOH)3 / 3 mol of NaOH)

= 4.10 x 10-4 mol of C3H5O(COOH)3 are titrated.

3) Calculating the mass of C3H5O(COOH)3 present in the juice sample.

Converting calculated moles of C3H5O(COOH)3 to grams by using the molar mass of C3H5O(COOH)3.

We know, the molar mass of C3H5O(COOH)3 = 192.124 g/mol

Thus,

= 4.1 x 10-4 mol of C3H5O(COOH)3 x ( 192.124 g / 1 mol)

= 0.0788 g of C3H5O(COOH)3 in the juice sample.

4) calculating the mass of C3H5O(COOH)3 present in 1 mL of apple juice.

= 1 mL apple juice x (0.07878 g of C3H5O(COOH)3 / 20.0 mL apple juice)

= 0.00394 g of C3H5O(COOH)3 in 1 mL of apple juice.